Can you please check this for me. It just does not seem right.
solve for x
4x(x-8)-5x(x-7)=-4
my answer is x=-1,4
Thank you

Found 2 solutions by Alan3354, nerdybill:
Answer by Alan3354(69443)

Can you please check this for me. It just does not seem right.
solve for x
4x(x-8)-5x(x-7)=-4
my answer is x=-1,4
--------------
Sub for x, see if it's right.
4*(-7) - 5*(-8) =? -4
-28 + 40 = ? -4 it's not right.
------------
4x(x-8)-5x(x-7)=-4
16*(4-8) - 20*(-3) =? -4
-64 + 60 = -4
----------
4 is right, -1 is not
Email your work via the TY note.

Answer by nerdybill(7384)

4x(x-8)-5x(x-7)=-4
4x^2-32x-5x^2+35x=-4
-x^2-32x+35x=-4
-x^2+3x=-4
-x^2+3x+4=0
x^2-3x-4=0
(x-4)(x+1) = 0
x = {-1, 5}
.
So, yes, you're right!